2) Find the area bounded by the curve x = y² + 2y and the line x = 3.

1
Expert's answer
2022-05-26T07:54:01-0400
"y^2+2y=3"

"y^2+2y+1=4"

"(y+1)^2=4"

"y_1=-3, y_2=1"

"A=\displaystyle\int_{-3}^{1}(3-(y^2+2y))dy"





"=\big[3y-\dfrac{y^3}{3}-y^2\big]\begin{matrix}
 1 \\
 -3
\end{matrix}"





"=3-\dfrac{1}{3}-1+9-9+9=\dfrac{32}{3}({units}^2)"